OpenStudy (anonymous):

An equation is shown below: 4(2x - 5) + 15 = 11 Write the steps you will use to solve the equation and explain each step

OpenStudy (ashes.boom):

first distribute the 4 to the 2x and the -5

OpenStudy (amorfide):

first of all you can subtract 15 from both sides to get the constants on one side 4(2x-5)+15 -15=11-15 4(2x-5)=-4 now you would divide by the scale factor, in this case 4 so divide by 4 on both sides so you can get the expression on its own then rearrange to get x

OpenStudy (anonymous):

so x =2 ??

OpenStudy (ashes.boom):

the way i did it i just distributed the 4 and got 8x -20 + 15 = 11 then you add 15 to -20 8x - 5 = 11 then you add 5 to both sides 8x = 16 then you can divide by 8 x = 2

OpenStudy (ashes.boom):

so yes you're right :)

OpenStudy (anonymous):

thank you

OpenStudy (ashes.boom):

Your welcome